# Defining implication relation for classical logic

### v11

@@ 3.2 Logic IRL
Let $\{A,B,...\}$ be a countable (finite or countably infinite) set of propositional constants representing atomic or primitive propositions.
@@
The set of all formulas is countably infinite as a subset of the set of finite-length strings over the countable alphabet $\{A,B,...\}\cup\{\top,\bot,\neg,\wedge,\vee,\to,\leftrightarrow\}$; and the set of all schemas is also countably infinite as a subset of the set of finite-length strings over the countable alphabet $\{\phi_{1},\phi_{2},...\}\cup\{\top,\bot,\neg,\wedge,\vee,\to,\leftrightarrow\}$.
@@
  • In a proof, by definition, any member of its proof sequence must be either a theorem (or an assumed theorem) or a result from the same-layer premise(s) by a rule (or by an assumed rule), not allowing any other cases (cf. Subsection 3.1). It is important to pay attention to this, especially in nested conditional proofs.
  • Rule of conditional proof of the general form $(\Gamma\cup\{\phi\}\vdash\psi)\Rightarrow(\Gamma\vdash\phi\to\psi)$ does not hold for IRL: If so, from $\{\phi,\psi\}\vdash\phi$ it follows $\phi\vdash\psi\to\phi$, hence $\vdash\phi\to(\psi\to\phi)$, contrary to that $\nvdash_{IRL}\phi\to(\psi\to\phi)$ (cf. Subsection 5.3).
The logic IRL is consistent as a sub-system of classical logic (cf. Subsection 5.3).
3.3 Basic derived results
The results in this subsection include mainly derived rules and meta rules.
@@ 4     $\phi[\phi\mapsto\psi]\leftrightarrow\phi$ 3: Selected replacement
5     $\psi\leftrightarrow\phi$ 4: Result of replacement
@@ 7 $\phi\leftrightarrow\psi\Rightarrow\psi\leftrightarrow\phi$ 6: Definition of derivable rule
Note: The definitions of conditional proof and derivable rule can be used tacitly in a conditional proof as usual. For example, the last two lines in the above proof can be omitted.
@@ 3 $(\phi\leftrightarrow\psi)[\psi\mapsto\chi]$ 2,1: Replacement
4 $\phi\leftrightarrow\chi$ 3: Result of Replacement
Leibniz’s replacement
Rule 3.3.3 (Leibniz’s replacement).
@@ 1 $\phi\leftrightarrow\psi$ Premise
2 $\chi\leftrightarrow\chi$ Reflexivity
3 $(\chi\leftrightarrow\chi)[\phi\mapsto\psi]$ 1,2: Replacement
Theorems to rules
Meta Rule 3.3.1.
$(\vdash\phi\leftrightarrow\psi)\Rightarrow(\phi\Rightarrow\psi).$
@@ 4 $\chi\leftrightarrow\chi[\phi\mapsto\psi]$ 3: Selected replacement
1 $\vdash\phi\leftrightarrow\psi$ Premise (begin meta proof)
2     $\phi$ Premise (begin proof)
3     $\phi\leftrightarrow\psi$ 1 (assumed theorem)
4     $\phi[\phi\mapsto\psi]$ 3,2: Replacement
5     $\psi$ 4: Result of replacement (end proof)
6 $\phi\vdash\psi$ 2–5: Definition of conditional proof
7 $\phi\Rightarrow\psi$ 6: Definition of derivable rule (end meta proof)
By this meta rule, from axioms, for example, $\phi\wedge\psi\leftrightarrow\psi\wedge\phi$, $(\phi\wedge\psi)\wedge\chi\leftrightarrow\phi\wedge(\psi\wedge\chi)$ and $(\phi\to\psi)\leftrightarrow(\phi\wedge\psi\leftrightarrow\phi)$, we obtain rules $\phi\wedge\psi\Rightarrow\psi\wedge\phi$, $(\phi\wedge\psi)\wedge\chi\Rightarrow\phi\wedge(\psi\wedge\chi)$ and $\phi\to\psi\Rightarrow\phi\wedge\psi\leftrightarrow\phi$, respectively.
@@ 7 $\phi\Rightarrow\psi$ 6: Definition of derivable rule (end meta proof)
Any formula of the form $(\phi\circ\psi)\leftrightarrow(\psi\circ\phi)$ (where $\circ$ stands for $\leftrightarrow$, $\wedge$ or $\vee$) is a theorem, so $(\psi\circ\phi)\leftrightarrow(\phi\circ\psi)$ is also a theorem by symmetry of $\leftrightarrow$, hence $\eta(\phi\circ\psi)\leftrightarrow\eta(\psi\circ\phi)$ and $\eta(\psi\circ\phi)\leftrightarrow\eta(\phi\circ\psi)$ are theorems by Leibniz’s replacement, and from these we obtain the bi-inference rule $\eta(\phi\circ\psi)\Leftrightarrow\eta(\psi\circ\phi)$. Thus, the symmetry properties can be used tacitly in a proof, viewing $\eta(\phi\circ\psi)$ and $\eta(\psi\circ\phi)$ as the same thing.
@@ 3 $\phi\leftrightarrow\phi\wedge\psi$ 1: Definition of implication
4 $\phi[\phi\mapsto\phi\wedge\psi]$ 3,2: Replacement
5 $\phi\wedge\psi$ 4: Result of replacement
6 $\psi$ 5: Conjunction elimination
Properties of theorems
@@ 3 $\top$ 2: Conjunction elimination
Meta Rule 3.3.2 (Theorem like top).
$(\vdash\psi)\Rightarrow(\vdash\phi\to\psi)\Leftrightarrow(\vdash\phi\wedge\psi\leftrightarrow\phi).$
@@ 4 $\vdash\phi\to\psi$ 2–3: Conditional proof
The second part $(\vdash\phi\to\psi)\Leftrightarrow(\vdash\phi\wedge\psi\leftrightarrow\phi)$ follows from the definition of implication. ∎
@@
Rule 3.3.5 (Theorem equivalence).
@@ 2 $\vdash\psi$ Premise
3     $\phi$ 1 (assumed theorem)
4     $\psi$ 2 (assumed theorem)
5     $\phi\wedge\psi\leftrightarrow\psi$ 3: Theorem like top
6     $\phi\wedge\psi\leftrightarrow\phi$ 4: Theorem like top
7     $\phi\leftrightarrow\psi$ 5,6: Transitivity of $\leftrightarrow$
3 $\vdash\phi\leftrightarrow\psi$ 3–7: Unconditional proof
Meta Rule 3.3.3 (Theorem representation).
@@ 3     $\top$ Top is a theorem
4     $\top[\top\mapsto\phi]$ 2,3: Replacement
5     $\phi$ 4: Result of replacement
3 $\vdash\phi$ 2–5: Unconditional proof
@@ 5 $\vdash\phi\leftrightarrow\top$ 2–4: Unconditional proof
For example, theorem $\phi\vee\neg\phi\leftrightarrow\top$ can be equivalently represented by $\phi\vee\neg\phi$.
Equivalence of the three forms
With the rules of commutativity and associativity for conjunction, $\phi_{1}\wedge\cdots\wedge\phi_{n}$ is well defined for any $n\geq 1$, so that the rules of conjunction introduction and elimination can be generalized to $\{\phi_{1},...,\phi_{n}\}\Rightarrow\phi_{1}\wedge\cdots\wedge\phi_{n}$ and $\phi_{1}\wedge\cdots\wedge\phi_{n}\Rightarrow\phi_{i}$ respectively.
Meta Rule 3.3.4 (Rule of conditional proof with multiple premises).
Proof.
0 $\phi_{1}\wedge\cdots\wedge\phi_{n}$ Premise 1 $\phi_{1}$ 0: Conjunction elimination $\vdots$ n $\phi_{n}$ 0: Conjunction elimination $\psi_{1}$ 1–n: First proof-sequence member of $\{\phi_{1},...,\phi_{n}\}\vdash\psi$vdots $\psi$ 1–n: Last proof-sequence member of $\{\phi_{1},...,\phi_{n}\}\vdash\psi$
Lemma 3.3.1.
(Two meta rules)
($(\{\phi_{1},...,\phi_{n}\}\Rightarrow\psi)\Rightarrow(\{\phi_{1},...,\phi_{n}\}\vdash\psi),\;(\vdash\phi_{1}\wedge\cdots\wedge\phi_{n}\to\psi)\Rightarrow(\{\phi_{1},...,\phi_{n}\}\vdash\psi).$
Proof.
  1. 1.
    ($(\{\phi_{1},...,\phi_{n}\}\Rightarrow\psi)\Rightarrow(\{\phi_{1},...,\phi_{n}\}\vdash\psi)$:
    1 $\{\phi_{1},...,\phi_{n}\}\Rightarrow\psi$ Premise (meta proof) 2     $\phi_{1}$ Premise (proof)     $\vdots$ 3     $\phi_{n}$ Premise (proof) 4     $\psi$ 2…3: 1 (assumed rule) 5 $\{\phi_{1},...,\phi_{n}\}\vdash\psi$ 2–4: Definition of conditional proof
  2. 2.
    $(\vdash\phi_{1}\wedge\cdots\wedge\phi_{n}\to\psi)\Rightarrow(\{\phi_{1},...,\phi_{n}\}\vdash\psi)$:
    1 $\vdash\phi_{1}\wedge\cdots\wedge\phi_{n}\to\psi$ Premise (meta proof) 2     $\phi_{1}$ Premise (proof)     $\vdots$ 3     $\phi_{n}$ Premise (proof) 4     $\phi_{1}\wedge\cdots\wedge\phi_{n}$ 2…3: Conjunction introduction 5     $\phi_{1}\wedge\cdots\wedge\phi_{n}\to\psi$ 1 (assumed theorem) 6     $\psi$ 5,4: Modus ponens 7 $\{\phi_{1},...,\phi_{n}\}\vdash\psi$ 2–6: Definition of conditional proof
Meta Rule 3.3.5 (Equivalence of the three forms).
Proof.
It follows from the definition of derivable rule, Meta Rule 3.3.4, and Lemma 3.3.1. ∎
If we have any one of the three forms, we get the other two. We just need to find either an unconditional proof $\vdash\phi_{1}\wedge\cdots\wedge\phi_{n}\to\psi$ or a conditional proof $\{\phi_{1},...,\phi_{n}\}\vdash\psi$ to obtain both the theorem $\phi_{1}\wedge\cdots\wedge\phi_{n}\to\psi$ and the rule $\{\phi_{1},...,\phi_{n}\}\Rightarrow\psi$.
3.4 Some theorems of IRL
This subsection demonstrates proofs of some theorems of IRL.
Basic theorems
@@ 9 $\neg\neg\phi\leftrightarrow\phi$ 1–8: Transitivity of bi-implication
Note: Leibniz’s replacement may be used tacitly, e.g. “LEM: Leibniz’s replacement” can be written as just “LEM”.
@@ 2 $\phi\to\phi$ 1: Definition of implication
Theorem 3.4.7 (Antisymmetry of implication / Bi-implication introduction).
$(\phi\to\psi)\wedge(\psi\to\phi)\to(\phi\leftrightarrow\psi).$
Proof.
1 $\phi\to\psi$ Premise
2 $\psi\to\phi$ Premise
3 $\phi\wedge\psi\leftrightarrow\phi$ 1: Definition of implication
4 $\psi\wedge\phi\leftrightarrow\psi$ 2: Definition of implication
5 $\phi\leftrightarrow\psi$ 3,4: Transitivity of bi-implication
Theorem 3.4.8 (Transitivity of implication).
$(\phi\to\psi)\wedge(\psi\to\chi)\to(\phi\to\chi).$
@@ 4 $\psi\wedge\chi\leftrightarrow\psi$ 2: Definition of implication
5     $\leftrightarrow\phi\wedge\psi\wedge\chi$ 3: Replacement
6     $\leftrightarrow\phi\wedge\psi$ 4: Replacement
@@ 4 $\phi\wedge(\neg\phi\vee\psi)\leftrightarrow\phi\wedge\psi$ 1–3: Transitivity
Theorem 3.4.9 (Equivalent definitions of implication).
$(\phi\to\psi)$
$\leftrightarrow$ $(\phi\wedge\psi\leftrightarrow\phi)$
$\leftrightarrow$ $(\phi\vee\psi\leftrightarrow\psi)$
$\leftrightarrow$ $(\phi\wedge\neg\psi\leftrightarrow\bot)$
$\leftrightarrow$ ($(\neg\phi\vee\psi\leftrightarrow\top).$

### v12

@@ 3.2 Logic IRL
Suppose there is a countable (finite or countably infinite) set $K$ of atomic or primitive propositions. Let $A,B,C,...$ be primitive propositional variables ranging over $K$.
@@
[…]
@@
  • In a proof, by definition, any member of its proof sequence must be either a theorem (or a theorem indicated in an outer layer), or a result from a subset of the premises and previous members in its own layer by a rule (or by a rule indicated in an outer layer), no other cases are allowed. It is important to pay attention to this, especially for a nested conditional proof (cf. Subsection 3.1).
  • Rule of conditional proof of the general form $(\Gamma\cup\{\phi\}\vdash\psi)\Rightarrow(\Gamma\vdash\phi\to\psi)$, where $\Gamma$ is a set of formulas, does not hold for IRL: if it holds, then from $\{\phi,\psi\}\vdash\phi$ it follows $\phi\vdash\psi\to\phi$, hence $\vdash\phi\to(\psi\to\phi)$, contrary to that $\nvdash_{IRL}\phi\to(\psi\to\phi)$ (cf. Subsection 5.3).
  • The logic IRL is consistent as a sub-system of classical (propositional) logic (cf. Remark 5.2.2 and Subsection 5.3).
3.3 Elementary derived results
@@ 4     $\phi[\phi\mapsto\psi]\leftrightarrow\phi$ 3: Selected replacement
5     $\psi\leftrightarrow\phi$ 4: Result of the replacement
@@ 7 $\phi\leftrightarrow\psi\Rightarrow\psi\leftrightarrow\phi$ 6: Definition of derivable rule
Note: The definitions of conditional proof and derivable rule can be used tacitly in a proof as usual. For example, the last two lines in the above proof can be omitted.
@@ 3 $(\phi\leftrightarrow\psi)[\psi\mapsto\chi]$ 2,1: Replacement
4 $\phi\leftrightarrow\chi$ 3: Result of the replacement
Equivalence of bi-implication theorem and bi-inference rule
Meta Rule 3.3.1 (Bi-implication theorem to bi-inference rule).
$(\vdash\phi\leftrightarrow\psi)\Rightarrow(\phi\Leftrightarrow\psi).$
Proof.
1 $\vdash\phi\leftrightarrow\psi$ Premise (begin meta proof)
2     $\phi$ Premise (begin proof 1)
3     $\phi\leftrightarrow\psi$ 1 (outside assumed theorem)
4     $\phi[\phi\mapsto\psi]$ 3,2: Replacement
5     $\psi$ 4: Result of the replacement (end proof 1)
6 $\phi\vdash\psi$ 2–5: Definition of conditional proof
7 $\phi\Rightarrow\psi$ 6: Definition of derivable rule
8     $\psi$ Premise (begin proof 2)
9     $\phi\leftrightarrow\psi$ 1 (outside assumed theorem)
10     $\psi\leftrightarrow\phi$ 9: Symmetry of bi-implication
11     $\psi[\psi\mapsto\phi]$ 10,8: Replacement
12     $\phi$ 11: Result of the replacement (end proof 2)
13 $\psi\vdash\phi$ 8–12: Definition of conditional proof
14 $\psi\Rightarrow\phi$ 13: Definition of derivable rule
15 $\psi\Leftrightarrow\phi$ 7,14: Definition of bi-inference rule (end meta proof)
By this meta rule, from bi-implication axioms we obtain bi-inference rules directly. For example, from $\phi\wedge\psi\leftrightarrow\psi\wedge\phi$, $(\phi\wedge\psi)\wedge\chi\leftrightarrow\phi\wedge(\psi\wedge\chi)$ and $(\phi\to\psi)\leftrightarrow(\phi\wedge\psi\leftrightarrow\phi)$, we obtain rules $\phi\wedge\psi\Leftrightarrow\psi\wedge\phi$, $(\phi\wedge\psi)\wedge\chi\Leftrightarrow\phi\wedge(\psi\wedge\chi)$ and $\phi\to\psi\Leftrightarrow\phi\wedge\psi\leftrightarrow\phi$, respectively.
Rule 3.3.3 (Bi-implication introduction).
$\{\phi\to\psi,\psi\to\phi\}\Rightarrow\phi\leftrightarrow\psi.$
Proof.
1 $\phi\to\psi$ Premise
2 $\psi\to\phi$ Premise
3 $\phi\wedge\psi\leftrightarrow\phi$ 1: Definition of $\to$
4 $\phi\leftrightarrow\phi\wedge\psi$ 3: Symmetry of $\leftrightarrow$
5 $\psi\wedge\phi\leftrightarrow\psi$ 2: Definition of $\to$
6 $\phi\wedge\psi\leftrightarrow\psi\wedge\phi$ Commutativity
7 $\phi\wedge\psi\leftrightarrow\psi$ 6,5: Transitivity of $\leftrightarrow$
8 $\phi\leftrightarrow\psi$ 4,7: Transitivity of $\leftrightarrow$
Meta Rule 3.3.2 (Bi-inference rule to bi-implication theorem).
$(\phi\Leftrightarrow\psi)\Rightarrow(\vdash\phi\leftrightarrow\psi).$
@@ 1 $\phi\Leftrightarrow\psi$ Premise
2       $\phi$ Premise
3       $\psi$ 2: The rule assumed in 1
4     $\phi\to\psi$ 2–3: Conditional proof
5     $\psi\to\phi$ 1–4: Similarly
6     $\phi\leftrightarrow\psi$ 4,5: Bi-implication introduction
7 $\vdash\phi\leftrightarrow\psi$ 4–6: Unconditional proof
Corollary 3.3.1 (Equivalence of bi-implication theorem and bi-inference rule).
$(\vdash\phi\leftrightarrow\psi)\Leftrightarrow(\phi\Leftrightarrow\psi).$
@@ 7 $\vdash\phi\leftrightarrow\psi$ 4–6: Unconditional proof
It follows from Meta Rules 3.3.1 and 3.3.2 by the definition of bi-inference rule. ∎
Thus, we do not need to distinguish bi-implication theorem $\vdash\phi\leftrightarrow\psi$ and bi-inference rule $\phi\Leftrightarrow\psi$: if we have one of the two, we have both.
@@
Meta Rule 3.3.3.
Let $F(\phi\circ\phi)$ be a schema containing $\phi\circ\psi$, where $\circ$ stands for $\leftrightarrow$, $\wedge$, or $\vee$. Then
$F(\phi\circ\psi)\Leftrightarrow F(\psi\circ\phi).$
Proof.
We already have $(\phi\circ\psi)\leftrightarrow(\psi\circ\phi)$ for $\leftrightarrow$, $\wedge$ and $\vee$, so the proof can be done below:
1     $F(\phi\circ\psi)$ Premise 2     $(\phi\circ\psi)\leftrightarrow(\psi\circ\phi)$ Known theorem 3     $F(\psi\circ\phi)$ 2,1: Replacement 4 $F(\phi\circ\psi)\Rightarrow F(\psi\circ\phi)$ 1–3: Definition of derivable rule 5 $F(\psi\circ\phi)\Rightarrow F(\phi\circ\psi)$ 1–4: Similarly 6 $F(\psi\circ\phi)\Leftrightarrow F(\phi\circ\psi)$ 4,5: Definition of bi-inference rule.
Thus, the symmetry properties of $\leftrightarrow$, $\wedge$ and $\vee$ can be used tacitly in a proof, viewing $F(\phi\circ\psi)$ and $F(\psi\circ\phi)$ as the same thing.
@@ 3 $\phi\leftrightarrow\phi\wedge\psi$ 1: Definition of implication
4 $\phi\wedge\psi$ 3,2: Replacement
5 $\psi$ 4: Conjunction elimination
Equivalence of the three forms
With the rules of commutativity and associativity for conjunction, $\phi_{1}\wedge\cdots\wedge\phi_{n}$ is well defined for any $n\geq 1$. And from the bidirectional reflexivity rule $\phi\Leftrightarrow\phi$ we get its unidirectional form $\phi\Rightarrow\phi$ by the definition of bi-inference rule. Therefore, the rules of conjunction introduction and elimination can be generalized to $\{\phi_{1},...,\phi_{n}\}\Rightarrow\phi_{1}\wedge\cdots\wedge\phi_{n}$ and $\phi_{1}\wedge\cdots\wedge\phi_{n}\Rightarrow\phi_{i}$ respectively, for all finite $n\geq 1$.
Meta Rule 3.3.4 (Rule of conditional proof with multiple premises).
Proof.
0 $\phi_{1}\wedge\cdots\wedge\phi_{n}$ Premise 1 $\phi_{1}$ 0: Conjunction elimination $\vdots$ n $\phi_{n}$ 0: Conjunction elimination $\psi_{1}$ 1…n: First proof-sequence member of $\{\phi_{1},...,\phi_{n}\}\vdash\psi$vdots $\psi$ 1…n: Last proof-sequence member of $\{\phi_{1},...,\phi_{n}\}\vdash\psi$
Meta Rule 3.3.5.
(Converse of CP rule)
$(\vdash\phi_{1}\wedge\cdots\wedge\phi_{n}\to\psi)\Rightarrow(\{\phi_{1},...,\phi_{n}\}\vdash\psi).$
Proof.
1 $\vdash\phi_{1}\wedge\cdots\wedge\phi_{n}\to\psi$ Premise (meta proof)
2     $\phi_{1}$ Premise (proof)
    $\vdots$
3     $\phi_{n}$ Premise (proof)
4     $\phi_{1}\wedge\cdots\wedge\phi_{n}$ 2…3: Conjunction introduction
5     $\phi_{1}\wedge\cdots\wedge\phi_{n}\to\psi$ 1 (assumed theorem)
6     $\psi$ 5,4: Modus ponens
7 $\{\phi_{1},...,\phi_{n}\}\vdash\psi$ 2–6: Definition of conditional proof
Meta Rule 3.3.6 (Equivalence of the three forms).
Proof.
It follows from the definition of derivable rule, Meta Rule 3.3.4 and Meta Rule 3.3.5, and the definition of bi-inference rule. ∎
Therefore, if we have any one of the three forms, we have the other two. So we just need to find either an unconditional proof $\vdash\phi_{1}\wedge\cdots\wedge\phi_{n}\to\psi$ or a conditional proof $\{\phi_{1},...,\phi_{n}\}\vdash\psi$ to obtain both the theorem $\phi_{1}\wedge\cdots\wedge\phi_{n}\to\psi$ and the rule $\{\phi_{1},...,\phi_{n}\}\Rightarrow\psi$.
Other elementary results
Rule 3.3.5 (Leibniz’s replacement).
Proof.
1 $\phi\leftrightarrow\psi$ Premise
2 $\chi\leftrightarrow\chi$ Reflexivity
3 $(\chi\leftrightarrow\chi)[\phi\mapsto\psi]$ 1,2: Replacement
@@ 3 $\top$ 2: Conjunction elimination
Meta Rule 3.3.7 (Theorem like top).
$(\vdash\psi)\Rightarrow(\vdash\phi\to\psi).$
@@ 4 $\vdash\phi\to\psi$ 2–3: Conditional proof
Meta Rule 3.3.8 (Theorem equivalence).
@@ 2 $\vdash\psi$ Premise
3 $\vdash\psi\to\phi$ 1: Theorem like top
4 $\vdash\phi\to\psi$ 2: Theorem like top
5     $\psi\to\phi$ 3 (theorem from an outer layer)
6     $\phi\to\psi$ 4 (theorem from an outer layer)
7     $\phi\leftrightarrow\psi$ 5,6: Bi-implication introduction
8 $\vdash\phi\leftrightarrow\psi$ 5–7: Unconditional proof
Meta Rule 3.3.9 (Theorem representation).
@@ 3     $\top$ Top is a theorem
4     $\phi$ 2,3: Replacement $[\top\mapsto\phi]$
5 $\vdash\phi$ 2–4: Unconditional proof
@@ 5 $\vdash\phi\leftrightarrow\top$ 2–4: Unconditional proof
For example, $\phi\vee\neg\phi\leftrightarrow\top$ and $\phi\vee\neg\phi$ are equivalent representations of LEM as a theorem.
3.4 More derivations in IRL
Some basic theorems or rules
Note that no matter what is proved is a theorem or a rule, we have both a theorem and a rule according to the equivalence of the three forms (Meta Rule 3.3.6) and the equivalence of bi-implication theorem and bi-inference rule (Corollary 3.3.1).
@@ 9 $\neg\neg\phi\leftrightarrow\phi$ 1–8: Transitivity of bi-implication
Note: Leibniz’s replacement may be used tacitly, e.g. “LEM: Leibniz’s replacement” can be written simply as “LEM” in a reason of a proof line.
@@ 2 $\phi\to\phi$ 1: Definition of implication
Antisymmetry of implication is the same as Bi-implication introduction (Rule 3.3.3).
Rule 3.4.1 (Transitivity of implication).
$\{\phi\to\psi,\psi\to\chi\}\Rightarrow\phi\to\chi.$
@@ 4 $\psi\wedge\chi\leftrightarrow\psi$ 2: Definition of implication
5     $\leftrightarrow\phi\wedge\psi\wedge\chi$ 3: Replacement $[\phi\mapsto\phi\wedge\psi]$
6     $\leftrightarrow\phi\wedge\psi$ 4: Replacement $[\psi\wedge\chi\mapsto\psi]$
@@ 4 $\phi\wedge(\neg\phi\vee\psi)\leftrightarrow\phi\wedge\psi$ 1–3: Transitivity
Rule 3.4.2 (Equivalent definitions of implication).
$\phi\to\psi$
$\Leftrightarrow$ $\phi\wedge\psi\leftrightarrow\phi$
$\Leftrightarrow$ $\phi\vee\psi\leftrightarrow\psi$
$\Leftrightarrow$ $\phi\wedge\neg\psi\leftrightarrow\bot$
$\Leftrightarrow$ $\neg\phi\vee\psi\leftrightarrow\top.$